OpenStudy (anonymous):

Let f(x) = 2x - 6 and g(x) = -3x - 7. Find (f(g))(3). I got 2 different answers, 26 and -38, i am not sure which one is right...

OpenStudy (lgbasallote):

find g(3) first g(3) = -3(3) - 7 g(3) = -9 - 7 g(3) = -16 sub this to every x in f(x) 2(-16) - 6 -32 - 6 -38 is the answer

OpenStudy (lgbasallote):

do you get how i did it?

OpenStudy (lgbasallote):

f(g(x))(3) = f(g(3))

OpenStudy (anonymous):

yes! i get it, my mistake was positive 32 instead of -32. Thanks!

OpenStudy (anonymous):

Find S10 for 1 + 4 + 7 +